# Federated Learning Assisted Edge Caching Scheme Based on Lightweight Architecture DDPM

### v2

@@ I-A Background
In recent years, the surge of smart devices has led to a significant increase in mobile data traffic, putting enormous pressure on wireless networks [1, 2, 3, 4]. As users become increasingly reliant on user devices such as smartphones and computers to access content, ensuring satisfactory service quality has become challenging [5, 6, 7, 8, 9]. To address this challenge, edge caching has become an effective solution [10, 11, 12, 13, 14, 15]. By caching user interested content in advance at wireless network edge nodes such as base stations (BS), user can directly obtain requested content from nearby BS instead of remote cloud server. This method can significantly alleviate network congestion, reduce traffic load, reduce service latency, and improve overall system performance [16, 17].
However, BSs have limited cache capacity, making it crucial to enhance the cache hit percentage of cached content [18, 19, 20]. Denoising diffusion probabilistic models (DPMs) have garnered significant attention for their superior generative capabilities, which generate samples through a step-by-step denoising process [21, 22, 23, 24]. Compared to earlier mainstream generative models (e.g., generative adversarial networks (GANs) and Flow-based models), DPMs exhibit more stable training processes and higher sample fidelity. However, the training of DPMs requires significant computational resources and is not suitable for resource-constrained devices. Recently, several lightweight architecture DPMs (LDPMs) have been proposed to address this challenge. In [25], Li et al. achieved sub-second text-to-image generation on mobile devices for the first time through the design of an efficient U-Net architecture and improved step distillation techniques. In [26], Chen et al. propose a lightweight DPM suitable for edge devices through a lightweight U-Net architecture design, which only requires 4 denoising steps to generate high-quality speech.
Additionally, model training requires access to users’ personal data. User personal data often contains a lot of privacy sensitive information, and users are unwilling to directly share their data with others, making it difficult to collect and train directly on user data [27, 28, 29, 30]. Fortunately, federated learning (FL) can address this issue by enabling the sharing of local models instead of raw user data [31, 32, 33, 34]. Therefore, it is necessary to introduce FL into edge caching in order to protect user privacy [35, 36, 37].
@@ I-B Related Work
Currently, there are many studies adopting FL in edge caching. In [38], Wang et al. combined FL with deep reinforcement learning to achieve decentralized collaborative caching, solving the communication bottleneck of centralized solutions. In [39], Yu et al. conducted edge caching for vehicle environments, integrating mobility prediction with FL to address the issues of high vehicle dynamics and user privacy. However, the aforementioned papers cannot accurately predict the content that users are interested in. The authors in [40] adopted a “K-nearest neighbor selection mechanism,” calculating similarity based on the interests between users, and using the interest lists of the K most similar “neighbor” users to the target user as auxiliary predictions of the target user’s interests. Although it can predict the content of user interest more accurately, predicting the content of interest to the target user requires the use of the “neighbor” user’s interest list, which to some extent exposes the privacy of the “neighbor” user. In [41], Wang et al. integrated Wasserstein GAN (WGANs) into FL frameworks, enabling accurate prediction without exposing raw user data. However, the training of GAN models demands substantial computational resources, rendering them impractical for deployment on resource-constrained user edge devices.
@@ II-B1 Forward Diffusion Process
In this work, we employ a varying scheduling strategy, where $\{\beta_{t}\}_{t=1}^{T}$ increases linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$. We define $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$, and can obtain $q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)$ and $\mathbf{x}_{t}$ as
@@ II-B2 Reverse Diffusion Process
Therefore, a simplified optimization function can obtain,
@@ III Edge Caching Scheme
Additionally, user data is typically high-dimensional and sparse, causing the Euclidean distances between data points to become uniform, and the noise distribution to become extremely flat [42, 43, 44, 45, 46]. This makes it difficult for the model to distinguish between signal and noise, and the DPM fails to effectively learn the distribution of user data [47, 48, 49, 50, 51, 52]. Therefore, we employ pre-trained encoder and decoder to process the data. Before performing local model training, we use the pre-trained encoder to map the raw user data into a low-dimensional latent space, allowing LDPM to learn the user data distribution in this low-dimensional space. Subsequently, when predicting content popularity at the BS, the pre-trained decoder is used to reconstruct the LDPM output data samples back to the original space dimensions. The pre-trained encoder and decoder are trained on the BS using publicly available datasets. Fig. 2 illustrates the framework of the encoder, decoder, and LDPM.
@@ III-A Federated-Based LDPM Training Algorithm
During the training process of FL, a total of $R_{\text{max}}$ rounds of training are conducted [53, 54, 55, 56]. Each round of training $r=1,2,\ldots,R_{\text{max}}$ consists of following four steps, corresponding to steps $1-4$ in Fig. 1.
@@ IV Simulation
Our overall U-Net architecture is similar to [57]. The main differences are the replacement of 2D convolutions with 1D convolutions to adapt to the user’s interaction data structure, and the use of one-fourth of the number of channels and three feature map resolutions to reduce the model size [58, 59, 60, 61]. Therefore, our LDPM has only 770K parameters, making it suitable for resource-constrained edge devices. To evaluate the scheme, we adopt the cache hit percentage and request content delay as evaluation metrics [62, 63, 64, 65, 66, 67]. The cache hit percentage represents the success rate of directly requesting content from the BS. The more accurate the predicted popular content, the higher the cache hit percentage. When the requested content is stored in the BS, it is regarded as a successful cache; conversely, if the content is not cached in the BS, it is termed a failed cache. The request content delay represents the average delay of all users getting content.
@@
  • Oracle : The Oracle algorithm possesses full prior knowledge of users future requests, defining the theoretical maximum achievable cache hit percentage.

### v3

@@ I-A Background
In recent years, the surge of smart devices has led to a significant increase in mobile data traffic, putting enormous pressure on wireless networks. As users become increasingly reliant on user devices such as smartphones and computers to access content, ensuring satisfactory service quality has become challenging [1]. To address this challenge, edge caching has become an effective solution. By caching user interested content in advance at wireless network edge nodes such as base stations (BS), user can directly obtain requested content from nearby BS instead of remote cloud server. This method can significantly alleviate network congestion, reduce traffic load, reduce service latency, and improve overall system performance [2].
However, BSs have limited cache capacity, making it crucial to enhance the cache hit percentage of cached content [3]. Denoising diffusion probabilistic models (DPMs) have garnered significant attention for their superior generative capabilities, which generate samples through a step-by-step denoising process [4]. Compared to earlier mainstream generative models (e.g., generative adversarial networks (GANs) and Flow-based models), DPMs exhibit more stable training processes and higher sample fidelity. However, the training of DPMs requires significant computational resources and is not suitable for resource-constrained devices. Recently, several lightweight architecture DPMs (LDPMs) have been proposed to address this challenge. In [5], Li et al. achieved sub-second text-to-image generation on mobile devices for the first time through the design of an efficient U-Net architecture and improved step distillation techniques. In [6], Chen et al. propose a lightweight DPM suitable for edge devices through a lightweight U-Net architecture design, which only requires 4 denoising steps to generate high-quality speech.
Additionally, model training requires access to users’ personal data. User personal data often contains a lot of privacy sensitive information, and users are unwilling to directly share their data with others, making it difficult to collect and train directly on user data [7]. Fortunately, federated learning (FL) can address this issue by enabling the sharing of local models instead of raw user data [8]. Therefore, it is necessary to introduce FL into edge caching in order to protect user privacy.
@@ I-B Related Work
Currently, there are many studies adopting FL in edge caching. In [9], Wang et al. combined FL with deep reinforcement learning to achieve decentralized collaborative caching, solving the communication bottleneck of centralized solutions. In [10], Yu et al. conducted edge caching for vehicle environments, integrating mobility prediction with FL to address the issues of high vehicle dynamics and user privacy. However, the aforementioned papers cannot accurately predict the content that users are interested in. The authors in [7, 11, 12] adopted a “K-nearest neighbor selection mechanism,” calculating similarity based on the interests between users, and using the interest lists of the K most similar “neighbor” users to the target user as auxiliary predictions of the target user’s interests. Although it can predict the content of user interest more accurately, predicting the content of interest to the target user requires the use of the “neighbor” user’s interest list, which to some extent exposes the privacy of the “neighbor” user. In [13], Wang et al. integrated Wasserstein GAN (WGANs) into FL frameworks, enabling accurate prediction without exposing raw user data. However, the training of GAN models demands substantial computational resources, rendering them impractical for deployment on resource-constrained user edge devices.
@@ II-B1 Forward Diffusion Process
In this work, we employ a varying scheduling strategy, where $\{\beta_{t}\}_{t=1}^{T}$ increases linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$. Similar to [14], we define $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$, and can obtain $q\left(\mathbf{x}_{t}|\mathbf{x}_{0}\right)$ and $\mathbf{x}_{t}$ as
@@ II-B2 Reverse Diffusion Process
In [4], Ho et al. proposed a simplified objective function for optimization, expressed as
@@ III Edge Caching Scheme
Additionally, user data is typically high-dimensional and sparse, causing the Euclidean distances between data points to become uniform, and the noise distribution to become extremely flat [11]. This makes it difficult for the model to distinguish between signal and noise, and the DPM fails to effectively learn the distribution of user data [15]. Therefore, we employ pre-trained encoder and decoder to process the data. Before performing local model training, we use the pre-trained encoder to map the raw user data into a low-dimensional latent space, allowing LDPM to learn the user data distribution in this low-dimensional space. Subsequently, when predicting content popularity at the BS, the pre-trained decoder is used to reconstruct the LDPM output data samples back to the original space dimensions. The pre-trained encoder and decoder are trained on the BS using publicly available datasets. Fig. 2 illustrates the framework of the encoder, decoder, and LDPM.
@@ III-A Federated-Based LDPM Training Algorithm
During the training process of FL, a total of $R_{\text{max}}$ rounds of training are conducted [8]. Each round of training $r=1,2,\ldots,R_{\text{max}}$ consists of following four steps, corresponding to steps $1-4$ in Fig. 1.
@@ IV Simulation
Our overall U-Net architecture is similar to [16]. The main differences are the replacement of 2D convolutions with 1D convolutions to adapt to the user’s interaction data structure, and the use of one-fourth of the number of channels and three feature map resolutions to reduce the model size [17, 18]. Therefore, our LDPM has only 770K parameters, making it suitable for resource-constrained edge devices. To evaluate the scheme, we adopt the cache hit percentage and request content delay as evaluation metrics [3]. The cache hit percentage represents the success rate of directly requesting content from the BS. The more accurate the predicted popular content, the higher the cache hit percentage. When the requested content is stored in the BS, it is regarded as a successful cache; conversely, if the content is not cached in the BS, it is termed a failed cache. The request content delay represents the average delay of all users getting content.
@@
  • Oracle [3]: The Oracle algorithm possesses full prior knowledge of users future requests, defining the theoretical maximum achievable cache hit percentage.